# fMoE: Fine-Grained Expert Offloading for Large Mixture-of-Experts Serving

### v1

@@ 3.1 Objectives and Challenges
How to avoid additional system overheads when managing experts? Our design must not introduce additional system overheads when serving existing MoE LLMs. We apply a series of system optimizations in fMoE to ensure serving efficiency and minimize additional overheads.
3.2 Architecture and Workflow
@@
Figure 5 describes the architecture and workflow of fMoE, which consists of three main components:
@@
Expert Map Store. We record expert maps, a new data structure defined in fMoE, to track fine-grained expert activation patterns from historical request prompts. expert maps provide nuance expert selection preferences over existing coarse-grained expert tracking methods (e.g., Expert Activation Matrix in MoE-Infinity [54]). The Expert Map Store dynamically keeps the most useful and unique expert maps for real-time queries during inference.
@@
Expert Map Matcher. When a request prompt arrives, fMoE searches the Expert Map Store for appropriate expert maps to guide expert prefetching before inference. expert map search is guided by calculating similarity scores in two folds: semantic and trajectory similarity.
@@
Expert Cache. After receiving the matched expert maps, fMoE prefetches experts from CPU memory to GPU memory for performing computations in inference. fMoE evicts and offloads low-priority expert weights to CPU memory if exceeding Expert Cache capacity.
@@
fMoE follows the five steps below to enable memory-efficient MoE serving with minimal inference latency:
@@
Step : Inference context collection. Before every inference iteration, fMoE collects necessary contexts, such as semantic embeddings and previous expert activation trajectories (§4.1), and feeds them to the Expert Map Matcher for hybrid similarity matching.
@@
Step : Expert map similarity matching. After receiving iteration-level contexts, the Expert Map Matcher finds and extracts the most similar expert maps by comparing the input context data with historical context data in the Expert Map Store (§4.2). The matched expert maps are forwarded to the Expert Cache to guide expert prefetching and offloading decisions.
@@
Step : Guided expert prefetching and offloading. We dynamically compute expert selection thresholds to determine which expert(s) to prefetch and offload in the MoE model guided by the searched expert maps (§4.3). Then, fMoE prefetches the expert weights from CPU to GPU memory and offloads cached experts from GPU to CPU when reaching the cache limit (§4.5).
@@
Step : Expert serving. The whole inference process consists of one iteration in the Prefill stage and multiple iterations in the Decode stage. For each MoE layer in every iteration, fMoE directly serves the expert required by the gating network if the corresponding weights are available in the GPU memory (defined as an expert hit). Otherwise, fMoE on-demand loads the expert weights from CPU to GPU to perform lossless serving (defined as an expert miss).
@@
Step : Expert map update. fMoE observes new expert maps produced after each iteration and updates them in the Expert Map Store (§4.4). When reaching the store capacity (e.g., 1K expert maps), fMoE deduplicates the Expert Map Store by identifying and dropping redundant expert maps to maintain diversity, maximizing the possibility of providing effective expert maps for any request prompts.
3.3 Problem Formulation
@@
We consider serving an MoE-based LLM with MoE layers on a GPU cluster, where each MoE layer has one gating network and experts. The gating network of each layer selects top experts for computation. The MoE model processes and generates answers for a workload consisting of unique request prompts. Each request prompt consists of multiple iterations processed during the prefill and decode stages, where is the request prompt collection. Let denote the -th expert at the -th layer in the -th iteration, where , , and . During each iteration , we can make at most prefetching decisions. Let and denote the set of cached experts and the set of activated experts for Iteration , respectively. Hence, we represent the result of whether an expert is hit (served by ) or miss (on-demand loading from CPU memory):
@@
where means is a miss. Since all experts in an MoE model are typically designed to have the same weight size, we assume experts’ loading time and memory footprint are homogenous.444We only consider selective experts. Some MoE models, such as Qwen1.5-MoE-A2.7B, have a few always-on experts that are not offloadable. Therefore, the total on-demand loading latency is summed across all iterations for each expert during the inference process, i.e., .
@@
The objective is to minimize the on-demand loading latency (ideally with perfect predictions) while limiting the total memory footprint of cached experts to satisfy the available GPU memory . Constraint 1 denotes the total number of prefetched experts should not exceed the total number of all experts in the MoE model. Constraint 2 represents the total number of activated experts, which must be the same as the total number of top experts summed across all layers. Constraint 3 describes the total memory footprint of prefetched experts must be limited by the available GPU memory size. Note that solving the ILP problem is already NP-hard [9], while in reality, prefetching experts always have mispredictions that further complicate the problem. Therefore, we opt for a heuristic-based design for fMoE.
4 fMoE’s Design
4.1 Expert Maps
We propose a new data structure, expert maps, to track expert activation patterns with a fine granularity. Figure 6 depicts the structure of an expert map. During the -th iteration, the -th self-attention layer first calculates the attention states. The gate network receives attentions and computes a probability distribution over all the experts at Layer :
@@ 4.1 Expert Maps
By tracking expert maps, we guide fMoE to discover fine-grained expert patterns—the iteration-level expert selection preferences via probability distributions. Intuitively, analyzing probability distributions enables fMoE to not only identify which experts are binarily selected or omitted, but also to assess the confidence or preference assigned to each expert from the perspective of the gate networks.
@@
The design of expert maps has two key advantages over existing coarse-grained expert tracking methods (e.g., MoE-Infinity [54] tracks the request-level expert hit counts). First, existing works only focus on aggregated request-level expert activations, whereas an expert map tracks individual iterations with detailed expert selections. Second, existing works only record the expert hit counts, whereas we track detailed probability distributions. Note that expert maps can easily recover coarse-grained information by applying a top selection operator to the probability distributions and aggregating expert counts over iterations, therefore generalizing to existing tracking methods. We evaluate fMoE against other tracking methods to show the effectiveness of expert maps in §6.5.
4.2 Expert Map Search
@@
When predicting and prefetching experts for MoE models, a prefetch distance is usually defined to avoid impacting inference latency [47]. Prefetch distance refers to the number of layers ahead that a prefetch instruction is issued before the target layer activates its experts, similar to the same term in memory prefetching [25]. Let denote the prefetch distance of the MoE model to serve. Figure 7 shows that fMoE employs two fine-grained search approaches to jointly match expert maps for guiding expert prefetching. Semantic search compares the input embeddings with historical embeddings to find expert maps with similar inputs, whereas trajectory search observes previous expert trajectories (i.e., probability distributions) and matches similar expert maps. We combine both semantic and trajectory features to improve fMoE’s map-matching and expert offloading accuracy. Two search approaches’ effectiveness is evaluated in §6.5.
Semantic-based expert map search. For the initial layers , due to the prefetch distance , existing solutions [16, 47, 54] cannot observe expert patterns for prediction and prefetching before the target layer is ready to activate experts. Thus, they usually define coarse-grained rules for prefetching initial layers. For example, MoE-Infinity [54] prefetches the most popular experts across all historical data points.
In contrast, fMoE leverages semantic hints from the input prompt to search for the most useful expert maps, requiring zero knowledge from the expert activation patterns. When serving request prompts and recording their expert maps, we record the semantic embeddings for each inference iteration: Since existing MoE-based LLMs all contain an embedding layer for analyzing the semantic meaning of inputs, it’s natural to extract the semantic embeddings using the output from the model’s original embedding layer. For any input prompts, we compute pairwise cosine similarity between the semantic embedding and the collection of historical semantic embeddings in the Expert Map Store:
@@
where is the batch size of input prompts, is the Expert Map Store capacity, and is the hidden dimension size. Then, for prompt , the historical Iteration with the highest score is selected. We use partial expert maps from the selected iteration, , to guide the expert prefetching for layers .
@@
Trajectory-based expert map search. Unlike layers , we can observe expert probability trajectories of previous layers to search expert maps for layers . Similar to the semantic-based search, we compute pairwise cosine similarity between the observed trajectory, , and the collection of historical expert maps, , in the Expert Map Store:
@@
We select the historical iteration with the highest score. Then, we use from the selected expert map to guide the expert prefetching for the target Layer , where is the prefetch distance.
@@
4.3 Expert Prefetching
Given the searched and customized expert map for a layer , we guide the expert prefetching in fine granularity.
Similarity-aware expert selection. With the different contexts collected during iterations, expert maps searched by fMoE also have varying similarity scores, which reflects the search confidence. To quantify the correlations between similarity score and expert hit rate, we run three MoE models (Mixtral-87B, Qwen1.5-MoE, and Phi-3.5-MoE) with two datasets (LMSYS-Chat-1M and ShareGPT) using the methodology described in §4.2. For each inference iteration, we compute the similarity scores, collect expert hit rates guided by the searched expert maps, and calculate the Pearson correlation coefficients [8]. Pearson coefficient is commonly used to measure correlations between variables, where a coefficient close to 1 indicates a strong positive correlation and a coefficient close to 0 means a weak correlation. Figure 8 shows the Pearson coefficient between similarity score and expert hit rate with three MoE models and two datasets. The results show that high similarity scores potentially relate to high expert hit rates if using the corresponding expert maps. Hence, we design fMoE’s expert prefetching to be similarity-aware.
For a layer with a to prefetch, we first dynamically compute an expert selection threshold given by
@@ 4.3 Expert Prefetching
where is the number of experts needed to activate per layer (e.g., Mixtral-87B activates two experts per layer). Constraint 7 requires the total probability of selected experts to prefetch per layer to be greater than . Constraint 8 represents the minimum amount of selected experts must be larger than the number of experts to activate required by the MoE model. Intuitively, we assign higher to low-score expert maps so that more experts are prefetched to mitigate mispredictions and assign lower for high-score expert maps to reduce memory footprint. Experts with higher probabilities are prioritized to be prefetched.
Asynchronous expert map matching and prefetching. Existing studies [16, 54] predict and prefetch experts synchronously during inference, severely hindering the inference performance. For example, MoE-Infinity [54] cannot compute forward functions before finishing expert prediction and prefetching at every MoE layer [55]. To minimize the system overhead and inference latency, we decouple the map matching and expert prefetching from the inference process using an asynchronous Publisher-Subscriber architecture (Figure 7). The Expert Map Store is a message broker that keeps messages from both the inference process and the Expert Map Matcher. As the inference proceeds, fMoE’s inference process continuously publishes and writes the inference contexts (i.e., semantic embeddings and expert probability distributions) to the Expert Map Store. At the same time, the Expert Map Matcher subscribes to the context data, matches expert maps based on new context data, and prefetches experts to the Expert Cache in an asynchronous manner.
4.4 Expert Map Store Management
Practically, we design fMoE’s Expert Map Store to maintain a capacity for storing unique expert maps. To effectively guide inference across diverse prompts, it makes sense to identify and deduplicate redundant expert maps.
Expert map deduplication. Since fMoE uses two approaches (i.e., semantic-based and trajectory-based) to compute similarity, we unify the two similarity scores to compute the pairwise redundancy scores between new iteration data and historical iteration data:

### v2

@@ 3.1. Objectives and Challenges
How to avoid additional system overheads when managing experts? Our design must not introduce additional system overheads when serving existing MoE LLMs. We apply a series of system optimizations in FineMoE to ensure serving efficiency and minimize additional overheads.
3.2. Architecture and Workflow
@@
Figure 5 describes the architecture and workflow of FineMoE, which consists of three main components:
@@
Expert Map Store. We record expert maps, a new data structure defined in FineMoE, to track fine-grained expert activation patterns from historical request prompts. expert maps provide nuance expert selection preferences over existing coarse-grained expert tracking methods (e.g., Expert Activation Matrix in MoE-Infinity (Xue et al., 2024)). The Expert Map Store dynamically keeps the most useful and unique expert maps for real-time inferences.
@@
Expert Map Searcher. When a request prompt arrives, FineMoE searches the Expert Map Store for appropriate expert maps to guide expert prefetching before inference. expert map search is guided by calculating similarity scores in two folds: semantic and trajectory similarity.
@@
Expert Cache. After receiving the searched expert maps, FineMoE prefetches experts from CPU memory to GPU to perform computations in inference. FineMoE evicts and offloads low-priority expert weights to CPU memory if exceeding Expert Cache capacity.
@@
FineMoE follows the five steps below to enable memory-efficient MoE serving with minimal inference latency:
@@
Step : Inference context collection. Before every inference iteration, FineMoE collects necessary contexts, such as semantic embeddings and previous expert activation trajectories (§4.1), and feeds them to the Expert Map Searcher for hybrid similarity searching.
@@
Step : Expert map similarity searching. After receiving iteration-level contexts, the Expert Map Searcher identifies the most similar expert maps by comparing the input context data with historical context data in the Expert Map Store (§4.2). The retrieved expert maps are forwarded to the Expert Cache to guide expert prefetching and offloading decisions.
@@
Step : Guided expert prefetching and offloading. We dynamically compute expert selection thresholds to determine which expert(s) to prefetch and offload in the MoE model guided by the searched expert maps (§4.3). Then, FineMoE prefetches the expert weights from CPU to GPU memory and offloads cached experts from GPU to CPU when reaching the cache limit (§4.5).
@@
Step : Expert serving. The whole inference process consists of one iteration in the Prefill stage and multiple iterations in the Decode stage. For each MoE layer in every iteration, FineMoE directly serves the expert required by the gating network if the corresponding weights are available in the GPU memory (defined as an expert hit). Otherwise, FineMoE on-demand loads the expert weights from CPU to GPU to perform lossless serving (defined as an expert miss).
@@
Step : Expert map update. FineMoE observes new expert maps produced after each iteration and updates them in the Expert Map Store (§4.4). When reaching the store capacity (e.g., 1K expert maps), FineMoE deduplicates the Expert Map Store by identifying and dropping redundant expert maps to maintain diversity, maximizing the possibility of providing effective expert maps for any request prompts.
3.3. Problem Formulation
@@
We consider serving an MoE-based LLM with MoE layers on a GPU cluster, where each MoE layer has one gating network and experts. The gating network of each layer selects top experts for computation. The MoE model processes and generates answers for a workload consisting of unique request prompts. Let denote the set of all requests, denote the set of all layers in a MoE model, and denote the set of all experts in a layer, respectively. Each request prompt consists of multiple iterations processed during the prefill and decode stages. Let denote the -th expert at the -th layer in the -th iteration, where , , and . During each iteration , we can make at most prefetching decisions. Let and denote the set of cached experts and the set of activated experts for Iteration , respectively. Hence, we represent the result of whether an expert is missed by :
@@
where means is a miss and requires on-demand loading from CPU memory. Since all experts in an MoE model are typically designed to have the same weight size, we assume experts’ loading time and memory footprint are homogenous.444We only consider selective experts. Some MoE models, such as Qwen1.5-MoE-A2.7B, have a few always-on experts that are not offloadable. Therefore, the total on-demand loading latency is summed across all iterations for each expert during the inference process:
@@
The objective is to minimize the on-demand loading latency (ideally with perfect predictions) while limiting the total memory footprint of cached experts to satisfy the available GPU memory . Constraint 1 denotes the total number of prefetched experts should not exceed the total number of all experts in the MoE model. Constraint 2 represents the total number of activated experts, which must be the same as the total number of top experts summed across all layers. Constraint 3 describes the total memory footprint of prefetched experts must be limited by the available GPU memory size. Note that solving the ILP problem is already NP-hard (Cormen et al., 2022), while in reality, prefetching experts always have mispredictions that further complicate the problem. Therefore, we opt for a heuristic-based design for FineMoE.
4. FineMoE’s Design
4.1. Expert Maps
We propose a new data structure, Expert Map, to track expert activation patterns with a fine granularity. Figure 6 depicts the structure of an expert map. During the -th iteration, the -th self-attention layer first calculates the attention states. The gate network receives attentions and computes a probability distribution over all the experts at Layer :
@@ 4.1. Expert Maps
By tracking expert maps, we guide FineMoE to discover fine-grained expert patterns—the iteration-level expert selection preferences via probability distributions. Intuitively, analyzing probability distributions enables FineMoE to not only identify which experts are binarily selected or omitted, but also to assess the confidence or preference assigned to each expert from the perspective of the gate networks.
@@
The design of expert maps has two key advantages over existing coarse-grained expert tracking methods (e.g., MoE-Infinity (Xue et al., 2024) tracks the request-level expert hit counts). First, existing works only focus on aggregated request-level expert activations, whereas an expert map tracks individual iterations with detailed expert selections. Second, existing works only record the expert hit counts, whereas we track detailed probability distributions. Note that expert maps can easily recover coarse-grained information by applying a top selection operator to the probability distributions and aggregating expert counts over iterations, therefore generalizing to existing tracking methods.
4.2. Expert Map Search
Given the historical expert maps defined in §4.1, FineMoE searches expert maps that provide the most accurate expert activation predictions with two fine-grained metrics: semantic similarity (§4.2.1) and trajectory similarity (§4.2.2). We also show that they are both effective in searching accurate historical expert maps for prediction and offloading (§4.2.3).
Existing solutions (Eliseev and Mazur, 2023; Song et al., 2024; Xue et al., 2024) cannot observe previous expert patterns for prediction and prefetching before the target layer is ready to activate experts for the initial layers , where represents the current layer index in an iteration and is referred as the prefetch distance. When predicting and prefetching experts for MoE models, prefetch distance is used to avoid impacting inference latency (Song et al., 2024; Zhang et al., 2025). Prefetch distance is the number of layers ahead that a prefetch instruction is issued before the target layer activates its experts, similar to the same term in memory prefetching (Lee et al., 2012). An ideal prefetch distance should perfectly overlap the prediction and prefetching operation overheads with the inference process.
Therefore, existing approaches (Eliseev and Mazur, 2023; Song et al., 2024; Xue et al., 2024) typically employ coarse-grained rules to prefetch experts for initial layers . For example, MoE-Infinity (Xue et al., 2024) prefetches the most popular experts across all historical data points. Even for layers , existing approaches use coarse-grained (request-level) metrics for predicting and prefetching experts, leading to low offloading accuracy.
@@ 4.2. Expert Map Search
In contrast, FineMoE leverages fine-grained iteration-level metrics tailored to the prefetch distance , employing semantic embeddings for layers prior to the prefetch distance and expert trajectories for layers subsequent to it. Figure 7 shows that FineMoE employs two fine-grained search approaches to jointly search expert maps for guiding expert prefetching: Semantic-based expert map search compares the input embeddings with historical embeddings to find expert maps with similar inputs, whereas trajectory-based search observes previous expert trajectories (i.e., probability distributions) and searches for similar expert maps. We combine both semantic and trajectory features to improve FineMoE’s map-searching and expert offloading accuracy.
4.2.1. Semantic-based Expert Map Search
Recent studies (Jie et al., 2025) demonstrate that semantic embeddings, i.e., embedding layer’s output after processing raw tokens, can potentially indicate expert selection behaviors. When serving request prompts and recording their expert maps, we record the semantic embeddings for each inference iteration. Existing MoE-based LLMs all contain an embedding layer for token semantic encoding, where words or subwords that appear in similar contexts will have similar embeddings (Mikolov et al., 2013). It’s natural to extract the semantic embeddings using the output from the model’s original embedding layer. Figure 7(a) shows the semantic-based expert map search in four steps: a1) extract semantic embeddings from the embedding layer, a2) compute similarity scores using semantic embeddings with historical data points in the Expert Map Store, a3) search similar expert maps based on similarity scores, and a4) prefetch experts with high probabilities for layers .
For any input prompts, we compute pairwise cosine similarity between the semantic embedding and the collection of historical semantic embeddings in the Expert Map Store:
@@ 4.2.1. Semantic-based Expert Map Search
where is the batch size of input prompts, is the Expert Map Store capacity, and is the hidden dimension size. Then, for prompt , the historical Iteration with the highest score is selected. We use partial expert maps from the selected iteration, , to guide layers .
4.2.2. Trajectory-based Expert Map Search
We leverage expert probability trajectories of previous layers to search expert maps for layers . Specifically, when , we use the past expert trajectories from Layer for prediction; when , we use the past trajectories from Layers and ; and so on. When (last layer), we use the past trajectories from Layers to for prediction. Figure 7(b) shows the trajectory-based expert search for a layer in four steps: b1) collect previous trajectory from Layers to , b2) compute similarity scores using collected trajectories with historical data points in the Expert Map Store, b3) search similar expert maps based on similarity scores, and b4) prefetch experts with high probabilities for the layer . We repeat this process until the last layer (Layer ) is completed.
@@ 4.2.2. Trajectory-based Expert Map Search
Similar to the semantic-based search, we compute pairwise cosine similarity between the observed trajectories, , and the collection of historical expert maps, , in the Expert Map Store:
@@
We select the historical iteration with the highest score. Then, we use from the selected expert map to guide the expert prefetching for the target layer .
@@
4.2.3. Effectiveness of Semantic and Trajectory Similarity
To verify how semantic and trajectory similarity scores can guide expert offloading, we run three MoE models (Mixtral-87B, Qwen1.5-MoE, and Phi-3.5-MoE) with two datasets (LMSYS-Chat-1M and ShareGPT). For each model and dataset, we first run prompts and record their semantic embeddings and expert trajectories, where each prompt generates one data point consisting of a semantic embedding and an expert map. Then, we exhaust all pairwise cases by calculating their semantic and trajectory similarity and expert hit rate (i.e., overlapped expert ratio). Figure 8 shows the mean expert hit rates of different semantic and trajectory similarity scores for three MoE models with LMSYS-Chat-1M. Both semantic and trajectory similarity can effectively indicate the accuracy of historical prompts or expert maps for offloading.
To statistically quantify the correlations between similarity score and expert hit rate, we calculate the Pearson correlation coefficients (Cohen et al., 2009) using all paired semantic and trajectory similarity scores and corresponding expert hit rates in Figure 8. The Pearson coefficient is commonly used to measure correlations between variables, where a coefficient close to 1 indicates a strong positive correlation and a coefficient close to 0 means a weak correlation. Figure 9 shows the Pearson coefficients between similarity score and expert hit rate with three MoE models and two datasets. The results show that high similarity scores potentially relate to high expert hit rates.
4.3. Expert Prefetching
Given the searched and customized expert map for a layer in Iteration , we explain how it guides FineMoE to dynamically prefetch experts in fine granularity.
Similarity-aware expert selection. With the different contexts collected during iterations, expert maps searched by FineMoE also have varying similarity scores.555In the following paper, we use “similarity scores” in both search contexts for simplicity, i.e., semantic similarity in semantic-based expert map search and trajectory similarity in trajectory-based search, respectively. Figures 8 and 9 demonstrated that similarity scores can effectively indicate the search confidence, where high searched similarity scores potentially mean high expert hit rates. Hence, we design FineMoE’s expert prefetching to be similarity-aware. For a layer with a to prefetch, we first dynamically compute an expert selection threshold :
@@ 4.3. Expert Prefetching
where is the number of experts needed to activate per layer (e.g., Mixtral-87B activates two experts per layer). Constraint 7 requires the total probability of selected experts to prefetch per layer to be greater than . Constraint 8 represents the minimum number of selected experts must be larger than the number of experts to activate required by the MoE model. Intuitively, we assign a higher to low-score expert maps so that more experts are prefetched to mitigate mispredictions and assign a lower for high-score expert maps to reduce the memory footprint. Experts with higher probabilities are prioritized to be prefetched.
Asynchronous expert map searching and prefetching. Existing studies (Eliseev and Mazur, 2023; Xue et al., 2024) predict and prefetch experts synchronously during inference, severely hindering the inference performance. For example, MoE-Infinity (Xue et al., 2024) cannot compute forward functions before finishing expert prediction and prefetching at every MoE layer (Xue, Leyang and Fu, Yao and Lu, Zhan and Mai, Luo and Marina, Mahesh, [n. d.]). To minimize the system overhead and inference latency, we decouple the map searching and expert prefetching from the inference process using an asynchronous Publisher-Subscriber architecture (Figure 7). The Expert Map Store is a message broker that keeps messages from both the inference process and the Expert Map Searcher. As the inference proceeds, FineMoE’s inference process continuously publishes and writes the inference contexts (i.e., semantic embeddings and expert probability distributions) to the Expert Map Store. At the same time, the Expert Map Searcher subscribes to the context data, searches expert maps based on new context data, and prefetches experts to the Expert Cache in an asynchronous manner.
4.4. Expert Map Store Management
Practically, we design FineMoE’s Expert Map Store to maintain a capacity for storing unique expert maps. To effectively guide inference across diverse prompts, it makes sense to identify and deduplicate redundant expert maps.
Expert map deduplication. Since FineMoE uses two approaches (i.e., semantic-based and trajectory-based) to compute similarity, we unify the two similarity scores to compute the pairwise redundancy scores between new iteration data and historical iteration data: